OpenStudy (danyjjb):

What will be the volume of ammonia at 30.0 C and 805 torr if its number is 21.0 L at 752 torr and 5.0 C

OpenStudy (sunnnystrong):

So you are going to need to use the combined gas law for this: P1*V1 / T1 = P2*V2 / T2 (805torr)(21.0L)/303.15K = (752torr)(V2)/278.15K Solve for V2 -->